# Generalized positive scalar curvature on spinc manifolds

### v2

@@ 3. Further Aspects of the “Generalized psc” Condition
Note that $\mathcal{R}^{+}(M)$ and $\mathcal{R}^{+,c}(M)$ are both open subsets of the separable Fréchet space of symmetric $2$-tensors on the tangent bundle, so they each have the homotopy type of a countable CW complex [20, Theorem 1]. We will show that the projection maps $\mathcal{R}^{+}(M)\to{\mathcal{C}^{+}(M)}$ and $p\colon\,\mathcal{R}^{+,c}(M)\to{\mathcal{C}^{+}(M)}$ are both fibrations with contractible fibers, and hence are homotopy equivalences. By the definitions, both of these maps are surjective, and $\mathcal{C}^{+}(M)$ has the quotient topology. Since $C^{\infty}_{+}(M)$ is a contractible Fréchet Lie group and it acts smoothly and freely on $\mathcal{R}^{+,c}(M)$, with $\mathcal{C}^{+}(M)$ the (paracompact) quotient space, the projection map $p\colon\,\mathcal{R}^{+,c}(M)\to\mathcal{C}^{+}(M)$ is a principal bundle for a contractible Fréchet Lie group, and thus is a homotopy equivalence. In particular, it has a section, which we can take to lie in $\mathcal{R}^{+}(M)$. We need to show that $\mathcal{R}^{+}(M)\to{\mathcal{C}^{+}(M)}$ is also a fibration with contractible fibers. We will show that the fibers are contractible open subsets of the fibers of $p$. Fix $C\in\mathcal{C}^{+}(M)$ and some metric $g_{0}\in p^{-1}(C)\cap\mathcal{R}^{+}(M)$. We will show that $p^{-1}(C)\cap\mathcal{R}^{+}(M)$ is convex, hence also contractible.
@@
Now we would like to do something similar for the curvature condition $R^{\text{{gen}}}_{(g,A)}>0$ on spinc manifolds. For this the paper [14], giving estimates on eigenvalues of the Dirac operator using the conformal Laplacian, should be replaced by [3, 13, 22].
@@ 6. The Index Difference Map
Now we turn to our spinc manifold $(M,L)$. We choose $(g,h,A)\in\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ and consider the spinc bordism $W=(M\times[0,1])\#N$ from $M$ to itself, where $N$ is attached to the interior of the cylinder $M\times[0,1]$. The bordism $W$ has natural spinc structure coming from $M$, as follows. Since $N$ is $2$-connected, the inclusion of either boundary component into $W$ is $2$-connected. Then we use a product metric $g+dt^{2}$ to extend the metric from $M\times\{0\}$ to $M\times[0,\varepsilon]$ for some $\varepsilon>0$, and use the pull-back of $L$ to $M\times[0,1]$ to get a spinc bundle on $W$. By the proof of the bordism theorem [5, Theorem 4.3] and the fact that the inclusion of either boundary copy of $M$ into $W$ is $2$-connected, we can push the triple $(g,h,A)$ across the bordism $W$ to obtain $(g^{\prime},h^{\prime},A^{\prime})\in\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ at $M\times\{1\}$. The value of the index difference is then obtained by computing the index of the Dirac operator on the manifold
@@
(6.1) $\sigma\colon\,\mathcal{R}^{+}(S^{k})\to\mathcal{R}^{{\mathrm{gen}+}}(S^{k}\times M)$
(up to homotopy) as follows. Given $g\in\mathcal{R}^{+}(S^{k})$, we can find a scale $t>0$ small enough so that the generalized scalar curvature
@@
is positive. Then the correspondence $g\mapsto(t^{2}g+g_{M},p^{*}A)$ defines $\sigma$ on any compact family $K\subset\mathcal{R}^{+}(S^{k})$ of metrics (for a suitable $t$ depending on $K$). In particular, the homomorphism $\sigma_{*}\colon\,\pi_{q}\mathcal{R}^{+}(S^{k})\to\pi_{q}\mathcal{R}^{{\mathrm{gen}+}}(S^{k}\times M)$ is well-defined. We obtain the following diagram:

### v3

@@ 3. Further Aspects of the “Generalized psc” Condition
Note that $\mathcal{R}^{+}(M)$ and $\mathcal{R}^{+,c}(M)$ are both open subsets of the separable Fréchet space of symmetric $2$-tensors on the tangent bundle, so they each have the homotopy type of a countable CW complex [20, Theorem 1]. We will show that the projection maps $\mathcal{R}^{+}(M)\to{\mathcal{C}^{+}(M)}$ and $p\colon\,\mathcal{R}^{+,c}(M)\to{\mathcal{C}^{+}(M)}$ are both fibrations with contractible fibers, and hence are homotopy equivalences. By the definitions, both of these maps are surjective, and $\mathcal{C}^{+}(M)$ has the quotient topology. Since $C^{\infty}_{+}(M)$ is a contractible Fréchet Lie group (under multiplication) and it acts smoothly and freely on $\mathcal{R}^{+,c}(M)$, with $\mathcal{C}^{+}(M)$ the (paracompact) quotient space, the projection map $p\colon\,\mathcal{R}^{+,c}(M)\to\mathcal{C}^{+}(M)$ is a principal bundle for a contractible Fréchet Lie group, and thus is a homotopy equivalence. In particular, it has a section. Choose such a section $s_{1}\colon\,\mathcal{C}^{+}(M)\to\mathcal{R}^{+,c}(M)$; it’s not canonical, but that won’t matter for us. So for each conformal class $C\in\mathcal{C}^{+}(M)$, $s_{1}(C)$ is a metric in the class $C$ for which $\mu_{1}(\mathsf{L}_{s_{1}(C)})>0$. There is a positive eigenfunction $\varphi_{C}$ for the eigenvalue $\mu_{1}(\mathsf{L}_{s_{1}(C)})$, and it’s unique up to multiplication by positive scalars. We can make it unique (on the nose!) by normalizing it to have maximum value $1$ on $M$. Then the map $C\mapsto\varphi_{C}$ is continuous in the appropriate topologies, because of elliptic regularity. Now we get a section $s\colon\,\mathcal{C}^{+}(M)\to\mathcal{R}^{+}(M)$ by taking $s(C)=\varphi_{C}^{\frac{4}{n-2}}s_{1}(C)$, using [18, Remark 2.12] to confirm that the scalar curvature of this metric is indeed strictly positive. We need to show that $\mathcal{R}^{+}(M)\to\mathcal{C}^{+}(M)$ is also a fibration with contractible fibers. We will show not only that the fibers are contractible open subsets of the fibers of $p$, but also that there is a deformation retraction from $\mathcal{R}^{+}(M)$ down to $s(\mathcal{C}^{+}(M))$ (a homeomorphic copy of $\mathcal{C}^{+}(M)$), and thus that $\mathcal{R}^{+}(M)$ and $\mathcal{C}^{+}(M)$ are homotopy equivalent.
Fix $C\in\mathcal{C}^{+}(M)$ and some metric $g_{0}\in p^{-1}(C)\cap\mathcal{R}^{+}(M)$. We will show that $p^{-1}(C)\cap\mathcal{R}^{+}(M)$ is convex in a suitable sense, hence also contractible.
@@
Now we get a deformation retraction of $\mathcal{R}^{+}(M)$ down to $s(\mathcal{C}^{+}(M))$ by using the above homotopy from a general metric $g_{1}\in\mathcal{R}^{+}(M)$ to $g_{0}=s(p(g_{1}))$, and thus $p$ restricted to $\mathcal{R}^{+}(M)$ is a homotopy equivalence. ∎
@@
Now we would like to do something similar for the curvature condition $R^{\text{{gen}}}_{(g,A)}>0$ on spinc manifolds (when $n\geq 3$). For this the paper [14], giving estimates on eigenvalues of the Dirac operator using the conformal Laplacian, should be replaced by [3, 13, 22]. What Bär calls the curvature endomorphism $\mathcal{K}$ is $\frac{1}{4}R_{g}+\frac{1}{2}ic(\Omega)=\frac{1}{4}R^{\text{{tw}}}_{(g,A)}$, which is a self-adjoint endomorphism at each point of $M$, and he takes $\kappa$ [3, bottom of p. 42] to be a scalar lower bound for $\mathcal{K}$, which since the eigenvalues of $ic(\Omega)$ come in $\pm$ pairs can be taken to be $\frac{1}{4}R^{\text{{gen}}}_{(g,A)}$, so the estimate from [3, Theorem 3] turns out to be exactly the inequality
$\lambda_{1}^{2}(D_{L})\geq\frac{n}{4(n-1)}\mu_{1}(\mathsf{L}_{(g,A)}),$
with $\mathsf{L}_{(g,A)}$ defined by replacing $R_{g}$ by $R^{\text{{gen}}}_{(g,A)}$ in the formula for $\mathsf{L}_{g}$, and this is nothing but the spinc analogue of Hijazi’s inequality.
@@ 6. The Index Difference Map
Now we turn to our spinc manifold $(M,L)$. We choose $(g,h,A)\in\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ and consider the spinc bordism $W=(M\times[0,1])\#N$ from $M$ to itself, where $N$ is attached to the interior of the cylinder $M\times[0,1]$. The bordism $W$ has natural spinc structure coming from $M$, since the spinc structure on $M$ defines one on the cylinder $M\times[0,1]$ and then this spinc structure extends uniquely to $W$ since $N$ is $2$-connected. We use a product metric $g+dt^{2}$ to extend the metric from $M\times\{0\}$ to $M\times[0,\varepsilon]$ for some $\varepsilon>0$. By the proof of the bordism theorem [5, Theorem 4.3] and the fact that the inclusion of either boundary copy of $M$ into $W$ is $2$-connected, we can push the triple $(g,h,A)$ across the bordism $W$ to obtain $(g^{\prime},h^{\prime},A^{\prime})\in\mathcal{R}^{{\mathrm{gen}+}}(M,L)$ at $M\times\{1\}$. The value of the index difference is then obtained by computing the index of the Dirac operator on the manifold
@@
(6.1) $\sigma\colon\,\pi_{q}(\mathcal{R}^{+}(S^{k}))\to\pi_{q}(\mathcal{R}^{{\mathrm{gen}+}}(S^{k}\times M))$
as follows. Fix a map $f\colon\,S^{q}\to\mathcal{R}^{+}(S^{k})$ representing a homotopy class, and note that we can find a scale $t>0$ small enough so that the generalized scalar curvature
@@
is positive for all $g\in f(S^{q})\subset\mathcal{R}^{+}(S^{k})$, the compact image of our map $S^{q}\to\mathcal{R}^{+}(S^{k})$. Then the formula $x\mapsto(t^{2}f(x)+g_{M},p^{*}A)$, $x\in S^{q}$, defines $\sigma([f])$, since it’s clear that up the homotopy class of this family of metrics only depends on the homotopy class of $f$ and not on the parameter $t$. Once $t$ is fixed, $\sigma$ is obtained by taking the product of the family of metrics $t^{2}f$ (which represents the same homotopy class as $f$) with the metric $g_{M}$ on $M$, and this operation of direct product is compatible with addition of homotopy classes, so in this way we get a homomorphism $\pi_{q}(\mathcal{R}^{+}(S^{k}))\to\pi_{q}(\mathcal{R}^{{\mathrm{gen}+}}(S^{k}\times M))$. We obtain the following diagram: